# Debt Behind the AI Boom: A Large-Scale Empirical Study of AI-Generated Code in the Wild

### 30 Mar 2026 v1

@@ 9. Conclusion
AI coding assistants are rapidly becoming part of real-world software development, but their long-term impact on software quality remains unclear. In this paper, we presented a large-scale empirical study of the technical debt introduced by AI-generated code in the wild. By mining 304,362 AI-authored commits from 6,275 GitHub repositories, we designed a commit-level pipeline to identify introduced technical debt and track its later evolution in production repositories. Our study provides a longitudinal view of how AI-generated code ages after it is merged, including what kinds of debt are introduced, how debt varies across assistants, and whether introduced issues persist or are later addressed. Overall, this work offers new evidence on the maintenance costs behind the rapid adoption of AI coding assistants and highlights the need for stronger quality assurance in AI-assisted software development.

### 26 Apr 2026 v2

@@ IX Conclusion and Future Work
AI coding assistants are rapidly becoming part of real-world software development, but their long-term impact on software quality remains unclear. In this paper, we presented a large-scale empirical study of the technical debt introduced by AI-generated code in the wild. By mining 302.6K AI-authored commits from 6,299 GitHub repositories, we designed a commit-level pipeline to identify introduced technical debt and track its later evolution in production repositories. Our study provides a longitudinal view of AI-generated code after it is merged. We look at what kinds of debt appear, how they differ across assistants, and whether they still remain at the latest revision. The results reveal the hidden maintenance costs behind AI coding assistants. They also show why stronger quality checks are needed in AI-assisted software development. In future work, we plan to extend our analysis to other forms of debt (e.g., architectural, documentation, and test-related debt), additional programming languages, and broader development contexts. We also plan to study what factors make AI-introduced debt more likely to persist. Finally, we hope our findings can inform the design of more debt-aware AI coding assistants that help developers catch and fix quality issues before they enter production.